x^2+8x+3 what are the x and y intersect?
Answer by Theo(13342)

it doesn't appear to factor easily so we'll use the completing the squares method.

x^2 + 8x + 3 becomes (x+4)^2 = -3 + 16 which becomes (x+4)^2 = 13

take the square root of both sides of that equation to get (x+4) = plus or minus sqrt(13)

subtract 4 from both sides of that equation to get x = -4 plus or minus sqrt(13)

it does appears to have x intecepts at x = -4 + sqrt(13) and x = - 4 - sqrt(13).

the y intercept will be the value of y when x is equal to 0.

y - x^2 + 8x + 3 has a y intercept at y = 3 because that's the value of y when x is equal to 0.

you should have:

y intercept at y = 3.
x intercepts at x = -4 + sqrt(13) and x = -4 - sqrt(13)

here's a graph of the equation.

note that sqrt(13) is roughly equal to 3.606.

x is therefore equal to -7.606 or -.394 when y is equal to 0.




$$$